OpenStudy (anonymous):

You buy a commemorative coin for $25. The value of the coin increases 3.25% per year. How much will the coin be worth in 15 years? Round to the nearest cent.

OpenStudy (anonymous):

so 25*3.25=81.25 81.25/100=.8125 .8125*15 years=12.18 12.18+25=37.18 im i right